Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250

4984 results found, displaying results 121 - 141.

Performance Issue: The CU Blog service is experiencing unstable service

Start:
2018-01-12 11:28:00
Status:
Closed
Brief Description:
The CU Blog service is currently experiencing intermittent service disruptions. We are actively investigating the issue with CampusPress.
Current Status:
Service has been restored. CampusPress is reviewing the issue to determine cause.
more details

Scheduled Service Change: Upgrade the bulkmail elist server (B1) to Aurea LM v12.1.21

Start:
2018-01-11 06:00:00
Status:
Closed
Brief Description:
Upgrade the bulkmail e-list server (B1) to the latest version of Aurea LM 12.1.21 (from v12.0m) to deploy bug fixes, security fixes and stay current with the product.
more details

Scheduled Service Change: Move DNSDB to New Host

Start:
2018-01-10 06:00:00
Status:
Closed
Brief Description:
Move the DNSDB Service to a new server.
more details

Performance Issue: Some reports of alumni not receiving messages sent from Cornell Office 365

Start:
2018-01-09 15:21:00
Status:
Closed
Brief Description:
Alumni may not be receiving some messages from Cornell Office 365 (the email service used by faculty and staff).
Current Status:
This issue is resolved. CIT is continuing to work with Microsoft to understand the cause and scope.
more details

Scheduled Service Change: Migrate Human Ecology to Media3 Production

Start:
2018-01-09 08:00:00
Status:
Closed
Brief Description:
Moving Human Ecology VMs from test to production at Media3
more details

Scheduled Service Change: Migrate Lightning pages and reports from UAT environment

Start:
2018-01-08 09:00:00
Status:
Closed
Brief Description:
Enable Lightning Experience & migrate pages/reports
more details

Scheduled Service Change: Upgrade the alternate elist server (Y2) to Aurea LM v12.1.21

Start:
2018-01-09 06:00:00
Status:
Closed
Brief Description:
Upgrade the alternate e-list server (Y2) to the latest version Aurea LM 12.1.21 (from v12.0m) to deploy bug fixes and stay current with the product.
more details

Scheduled Service Change: Emergency registry changes for Meltdown Spectre

Start:
2018-01-08 15:30:00
Status:
Closed
Brief Description:
Per Microsoft recommendation, deploy registry changes to address Spectre/Meltdown security issues across Managed Server Windows machines. A reboot will be required.
more details

Performance Issue: Different Login Screen When Logging Into Outlook

Start:
2017-12-19 12:03:00
Status:
Closed
Brief Description:
Due to a change in how authentication is handled by Microsoft services, some individuals using Outlook for Windows may see a screen they are unfamiliar with, asking them to log in.
Current Status:
This login will continue to be used when logging into Microsoft services.
more details

Performance Issue: Mail from E-List and Certain Other Cornell Sources to Office 365

Start:
2018-01-03 09:30:00
Status:
Closed
Brief Description:
At around 9:30am, Wednesday, January 3, 2018, Office 365 began rejecting email sent from e-list and some other Cornell systems. At 3:30pm Microsoft made a change that allowed the mail to begin being accepted again.
Current Status:
Email from all Cornell sources that were affected by this problem continues to be delivered to Office 365.
more details

Scheduled Service Change: Schurman Hall - Voice & Data Network Connectivity

Start:
2018-01-07 07:00:00
Status:
Closed
Brief Description:
Change Description: Second day rescheduled from 12/28/2017 to 1/7/2018. Re: CRQ000001022935 (see: https://itservicealerts.hosting.cornell.edu/view/5042)
more details

Scheduled Service Change: OS patching bulk mail e-list server (B1)

Start:
2018-01-07 07:45:00
Status:
Closed
Brief Description:
Change Description: The bulk mail e-list server (B1) will be down for OS patching
more details

Scheduled Service Change: Update certificate on LDS servers

Start:
2018-01-07 08:00:00
Status:
Closed
Brief Description:
Change Description: Update the expiring certificates on leeds26 and leeds26. The certs are due to expire on January 11.
more details

Scheduled Service Change: Routine Maintenance of User Edge Switches and Wireless Controllers

Start:
2018-01-07 05:00:00
Status:
Closed
Brief Description:
REMINDER: On Sunday, January 7, 2018, from 5:00am to 8:00am, CIT Network Engineering will perform routine maintenance on all user edge network switches and wireless controllers.
more details

Scheduled Service Change: TeamForge/SVN/SourceForge Prod Linux OS update cycle

Start:
2018-01-07 04:00:00
Status:
Closed
Brief Description:
Change Description: TeamForge/SVN/SourceForge Prod Linux OS update cycle
more details

Scheduled Service Change: Upgrade Core Routers

Start:
2018-01-07 05:00:00
Status:
Closed
Brief Description:
Upgrade core routers (one at a time) to the recommended maintenance release. All core services are redundant.
more details

Scheduled Service Change: Upgrade On-Premise Amazon Web Services (AWS) Direct Connect Routers

Start:
2018-01-07 05:00:00
Status:
Closed
Brief Description:
Upgrade On-Premise Amazon Web Services (AWS) Direct Connect Routers (Primary & Backup)
more details

Performance Issue: EH&S staff phones unavailable due to renovations

Start:
2018-01-02 09:24:00
Status:
Closed
Brief Description:
Due to renovations, many EH&S staff do not have an active phone. Please call the main EH&S number, 255-8200, or the individual's cell phone, for the rest of this week.
Current Status:
EH&S Customer Service is notifying all building coordinators.
more details

Unplanned Outage: Wired and Wi-Fi data network outage in Wilson Lab

Start:
2018-01-04 11:37:00
Status:
Closed
Brief Description:
The wilson2-3500 switch is currently down.
Current Status:
The wilson2-3500 switch is now back up and wired and Wi-Fi data service has been restored.
more details

Performance Issue: Remedy outgoing e-mail spam/spoofing warnings

Start:
2018-01-02 10:00:00
Status:
Closed
Brief Description:
Remedy outgoing e-mail spam/spoofing warnings is being added by Cornell’s e-mail service provider, "[This sender failed our fraud detection checks and may not be who they appear to be. Learn about spoofing at http://aka.ms/LearnAboutSpoofing]”.
Current Status:
Steps have been taken to address this issue for Remedy and Identity Management app emails. Documentation is being worked on to address this subject in detail.
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250